# AdaptEval: Evaluating Large Language Models on Domain Adaptation for Text Summarization

### v1

@@ Abstract
Florian Matthes Technical University of Munich matthes@tum.de                        Laura Mascarell Espuny ETH Zürich lmascarell@inf.ethz.ch
@@ 1 Introduction
The experimental results show the abilities of LLMs to adapt to the domain in the ICL setting. In particular, small models with 7b parameters achieve comparable performance to their larger counterparts with only two learning examples. However, G-eval highlights the difficulty of adapting to the medical domain. While the fine-tuned models achieve the best performance in terms of automatic scores, their adaptation to the domain vocabulary is inferior to the ICL setting.  Finally, we release the domain benchmark and evaluation metrics as the first domain Adaptation Evaluation suite (AdaptEval) to facilitate the evaluation of models and foster further research on this task. [^1] GitHub link will be available upon publication.
@@ 2.2 Evaluation Metrics
Mistral 7b 0.863 35.81 27.17 0.863 34.00 27.29 0.833 21.66 23.08
Table 3: BERTScore F1, DVO (%), and the geometric mean of ROUGE-1/2/L (ROUGE) of all models across the three domains. The value ‘x’ implies that the model was not evaluated under those settings. Llama2  7b1/2/3 are fine-tuned with 1k, 5k, and 10k instances, respectively. See complete table in Appendix B.
@@ 5 Conclusion
Supplementary Materials Statement:
Source code of the AdaptEval suite along with the fine-tuned models will be available on GitHub and HuggingFace upon publication, respectively.
@@ A.3 Fine-tuning Setting
We included Llama2 (7b) [^18] https://huggingface.co/meta-llama/Llama-2-7b, Llama2 (13b) [^19] https://huggingface.co/meta-llama/Llama-2-13b, and Mistral AI  [^20] https://huggingface.co/mistralai/Mistral-7B-v0.1 for LLM fine-tuning. We fine-tuned the models for 1 epoch using the HuggingFace Trainer API and LoRA on a training subset consisting of samples with a maximum length of 4096, such that they can fit in the context window without truncation. Since Zhou et al. (2023) argue that 1k samples are enough to fine-tune LLMs, we experimented with 1k, 5k, and 10k training samples. We selected the LoRA parameters r=64, alpha=16, and a dropout of 0.1. Furthermore, we used the paged AdamW optimizer with a beta2 value of 0.999 and a learning rate of $42e-$ 4 with a constant learning rate strategy. We did not fine-tune Vicuna, since we only used the non-instruction tuned models in this setting. We excluded Falcon from fine-tuning as it only supports a context window of 2048, and therefore, it cannot be fairly compared against the other models with a context window of 4096.
@@ Appendix B Automatic Scores and DVO
Llama2 13b 0.862 35.28 26.26 0.860 32.67 26.47 x x x
Mistral 7b 0.863 35.81 27.17 0.863 34.00 27.29 0.833 21.66 23.08
Table 6: BERTScore F1, DVO (%), and the geometric mean of ROUGE-1/2/L (ROUGE) of all models across the three domains. The value ‘x’ implies that the model was not evaluated under those settings. Llama2  7b1/2/3 are fine-tuned with 1k, 5k, and 10k instances, respectively.

### v2

@@ Abstract
Florian Matthes Technical University of Munich matthes@tum.de                        Laura Mascarell ETH Zürich lmascarell@inf.ethz.ch
@@ 1 Introduction
The experimental results show the abilities of LLMs to adapt to the domain in the ICL setting. In particular, small models with 7b parameters achieve comparable performance to their larger counterparts with only two learning examples. However, G-eval highlights the difficulty of adapting to the medical domain. While the fine-tuned models achieve the best performance in terms of automatic scores, their adaptation to the domain vocabulary is inferior to the ICL setting.  Finally, we release the domain benchmark and evaluation metrics as the first domain Adaptation Evaluation suite (AdaptEval) to facilitate the evaluation of models and foster further research on this task. [^1] AdaptEval code is available on AdaptEval.
@@ 2.2 Evaluation Metrics
Mistral 7b2 0.863 35.81 27.17 0.863 34.00 27.29 0.833 21.66 23.08
Table 3: BERTScore F1, DVO (%), and the geometric mean of ROUGE-1/2/L (ROUGE) of all models across the three domains. The value ‘x’ implies that the model was not evaluated under those settings. 1/2/3 indicate fine-tuning with 1k, 5k, and 10k instances, respectively. See complete table in Appendix B.
@@ A.3 Fine-tuning Setting
The links to all fine-tuned models is displayed in Table 6.
Science Medical Government
BART bart-arxiv-1024 bart-pubmed-1024 bart-govreport-1024
PEGASUS-X bigbird-pegasus-arxiv-4096 bigbird-pegasus-pubmed-4096 bigbird-pegasus-govreport-4096
Llama2  7b1 Llama-2-7b-arxiv-4096 Llama-2-7b-pubmed-4096 Llama-2-7b-govreport-4096
Llama2  7b2 Llama-2-7b-arxiv-4096 Llama-2-7b-pubmed-4096 x
Llama2  7b3 Llama-2-7b-arxiv-4096 Llama-2-7b-hf-pubmed-4096 x
Llama2 13b2 Llama-2-13b-arxiv-4096 Llama-2-13b-pubmed-4096 x
Mistral 7b2 Mistral-7B–arxiv-4096 Mistral-7B-pubmed-4096 Mistral-7B-govreport-4096
Table 6: Links to all fine-tuned models repositories. The value ‘x’ implies that the model was not evaluated under those settings. 1/2/3 indicate fine-tuning with 1k, 5k, and 10k instances, respectively.
@@
We included Llama2 (7b) [^18] https://huggingface.co/meta-llama/Llama-2-7b, Llama2 (13b) [^19] https://huggingface.co/meta-llama/Llama-2-13b, and Mistral AI  [^20] https://huggingface.co/mistralai/Mistral-7B-v0.1 for LLM fine-tuning. We fine-tuned the models for 1 epoch using the HuggingFace Trainer API and LoRA on a training subset consisting of samples with a maximum length of 4096, such that they can fit in the context window without truncation. Since Zhou et al. (2023) argue that 1k samples are enough to fine-tune LLMs, we experimented with 1k, 5k, and 10k training samples. Since models do not show any performance increase when trained on more than 5k samples, we opted to train on Llama2 (13b) and Mistral AI on 5k samples. We selected the LoRA parameters r=64, alpha=16, and a dropout of 0.1. Furthermore, we used the paged AdamW optimizer with a beta2 value of 0.999 and a learning rate of $42e-$ 4 with a constant learning rate strategy. We did not fine-tune Vicuna, since we only used the non-instruction tuned models in this setting. We excluded Falcon from fine-tuning as it only supports a context window of 2048, and therefore, it cannot be fairly compared against the other models with a context window of 4096.
@@ Appendix B Automatic Scores and DVO
Llama2 13b2 0.862 35.28 26.26 0.860 32.67 26.47 x x x
Mistral 7b2 0.863 35.81 27.17 0.863 34.00 27.29 0.833 21.66 23.08
Table 7: BERTScore F1, DVO (%), and the geometric mean of ROUGE-1/2/L (ROUGE) of all models across the three domains. The value ‘x’ implies that the model was not evaluated under those settings. 1/2/3 indicate fine-tuning with 1k, 5k, and 10k instances, respectively.